# Student Member, IEEE

### v1

@@ Abstract
Continuously optimizing sensor placement is essential for precise target localization in various military and civilian applications. While information theory has shown promise in optimizing sensor placement, many studies oversimplify sensor measurement models or neglect dynamic constraints of mobile sensors. To address these challenges, we employ a range measurement model that incorporates radar parameters and radar-target distance, coupled with Model Predictive Path Integral (MPPI) control to manage complex environmental obstacles and dynamic constraints. We compare the proposed approach against stationary radars or simplified range measurement models based on the Root Mean Squared Error (RMSE) of the Cubature Kalman Filter (CKF) estimator for the targets’ state. Additionally, we visualize the evolving geometry of radars and targets over time, highlighting areas of highest measurement information gain, demonstrating the strengths of the approach. The proposed strategy outperforms stationary radars and simplified range measurement models in target localization, achieving a 38-74% reduction in mean RMSE and a 33-79% reduction in the upper tail of the 90% Highest Density Interval (HDI) over 500 Monte Carlo (MC) trials across all time steps. Code will be made publicly available upon acceptance.
@@ 1 INTRODUCTION
A key challenge in radar-based target tracking is in incorporating realistic sensor models into the information metrics used for controlling the sensors. In particular, most existing control strategies for this problem use the Posterior Fisher Information Matrix (PFIM) [12] (inverse of the Posterior Cramer Rao Lower Bound (PCRLB)) to quantify the relationship between radar-target geometry and target localization performance [13, 14, 15]. However, studies optimizing sensor placement often oversimplify by using a constant covariance matrix for range measurements, neglecting signal path attenuation effects [4]. The approach in [4] focuses on optimizing the angle subtended between the radars and targets but does not encourage the mobile radars to move closer to the target. In [16], the authors provide the error covariance matrix of a Best Linear Unbiased Estimator (BLUE) estimator for the range measurement using a first-order Taylor Series approximation of a distance dependent range (DDR) measurement model, which yields the Fisher Information Matrix (FIM). Similarly, we calculate the FIM for the the DDR measurement model from [16], where the covariance matrix of the range measurement is directly proportional to the received power at the radar after signal path attenuation, but we do not apply a first-order Taylor Series approximation to the DDR measurement model. This strategy encourages mobile radars to move closer to the targets, thereby increasing the Signal Noise Ratio (SNR) and improving radar range measurements.
@@
Another major challenge in scenarios with moving targets and movable radars is that the optimal radar-target geometry may change substantially over time. Consequently, many studies use Model Predictive Controller (MPC) [17] for choosing the control inputs of mobile sensors over a finite horizon in response to uncertain target state estimates [18, 19, 20, 21, 22]. However, natural choices for the MPC objective in this problem, such as minimizing/maximizing the D-optimality of PCRLB/PFIM, lead to challenging nonlinear optimization problems that are computationally prohibitive to solve exactly. To address this issue, we instead utilize Model Predictive Path Integral (MPPI) to approximate the optimal control inputs for the mobile radars, enabling the handling of discontinuous, non-smooth objectives through highly parallelized computation for real-time control [23, 24]. However, as the number of agents, planning horizon, and the actuation space increases, parameter estimation, particularly for the covariance, requires an exponentially growing number of samples. To mitigate this issue and reduce the number of proposal distribution updates and trajectory samples needed in the high-dimensional control space, we integrate advanced Adaptive Importance Sampling (AIS) techniques, such as cross-entropy weighting with covariance shrinkage [25], to prevent overfitting of the sample covariance matrix.
@@
This paper introduces a improved pipeline for continuously optimizing radar placement, aiming to overcome the previously mentioned challenges. We derived objective functions using the FIMs that incorporate signal path attenuation and optimize these functions with respect to radar control inputs using MPPI control. Crucially, MPPI control provides increased flexibility for real-time optimization of nonlinear, discontinuous, and non-differentiable objectives in research, including future enhancements to signal measurement models and occupancy grid maps. Results from over 500 Monte Carlo (MC) trials across various simulation scenarios empirically show that our approach achieves lower Root Mean Squared Error (RMSE) compared to stationary radars and state estimators using simplified range measurement models, with reduced variance across MC trials.
@@ 2 Related Work
We explore literature using the FIM of range measurements’ likelihood to optimize sensor trajectories, with the likelihood parameters as the target positions. Various forms of the FIM establish a fundamental estimation lower bound limit for the mean squared error covariance matrix [26]. Therefore, many authors construct an objective as the D/A-optimality [27] of the PFIM, parameterized by sensor and estimated target positions [28]. We divide these works into two categories: those that consider static-sensors versus those that consider dynamic-sensors.
@@ 2.2 Dynamic Sensors
Similar to information based MPC objectives like the D/A optimality of the FIMs, [37] minimizes the mutual information gain between the predict and the update steps of a linear Kalman Filter, while parameterizing the noise covariance of the measurement based on the robot’s state. However, the measurement covariance only scales linearly with the sensor-target distance and is clipped at an extremely small variance of 0.01 meters, restricting experiments to small square arenas of 40 to 60 meters.
@@ 3 Preliminaries
    Perfect data association. We assume that the process of selecting which measurement(s) to incorporate into the target state estimator, is perfect, and the radars constantly track the targets.
@@ 4 Radar Signal Model
and the set of orthonormal lowpass equivalent signals $s(t)$ such that there is no signal interference when multiple radars transmit simultaneously. The return signal noise $\epsilon_{a}(t)$ is circularly symmetric, zero-mean, complex Gaussian noise with autocorrelation function $\sigma_{a}^{2}\delta(\tau)$. $\zeta$ is the complex amplitude of the return signal, where $|\zeta|^{2}$ is the radar power return specified by the radar equation [39]:
@@
We denote all the radar parameters, signal parameters, and noise parameters as a constant $\Gamma=\frac{c^{2}\sigma_{a}^{2}(4\pi)^{3}L}{P_{t}\Lambda_{t}\Lambda_{r}\lambda^{2}\Xi 8\pi^{2}f_{c}^{2}}$. For brevity, we define shorthand notation as $h^{nm}=2\lVert\bm{\chi}^{R_{n}}_{xyz}-\bm{\chi}^{T_{m}}_{xyz}\rVert$ and $\sigma_{z}^{nm}=\Gamma\lVert{\bm{\chi}^{R_{n}}_{xyz}-\bm{\chi}^{T_{m}}_{xyz}}\rVert^{4}$.
@@ 5.1 Target Kinematics: Constant Velocity Model
The target kinematic model for a single target follows constant velocity motion with acceleration noise [41], which specifies the single-target transition model:
$\bm{\chi}^{T}_{k+1}$ $=A\bm{\chi}_{k}^{T}+\epsilon_{w}$ (5)
@@
$A_{\text{single}}$ $=\begin{bmatrix}1&0&0&\Delta t&0&0\\ 0&1&0&0&\Delta t&0\\ 0&0&1&0&0&\Delta t\\ 0&0&0&1&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\\ \end{bmatrix}$ (6)
@@
$W_{\text{single}}$ $=\begin{bmatrix}\frac{\Delta t^{4}}{4}&0&0&\frac{\Delta t^{3}}{2}&0&0\\ 0&\frac{\Delta t^{4}}{4}&0&0&\frac{\Delta t^{3}}{2}&0\\ 0&0&\frac{\Delta t^{4}}{4}&0&0&\frac{\Delta t^{3}}{2}\\ \frac{\Delta t^{3}}{2}&0&0&\Delta t^{2}&0&0\\ 0&\frac{\Delta t^{3}}{2}&0&0&\Delta t^{2}&0\\ 0&0&\frac{\Delta t^{3}}{2}&0&0&\Delta t^{2}\\ \end{bmatrix}\cdot\sigma_{W}^{2}$ (7)
$\sigma_{W}$ determines the degree of deviation of the dynamics from constant velocity, determined by the strength of the acceleration variance. Thus, the targets’ transition model (for all targets concatenated as a column vector) is:
@@
$\mathbf{z}^{T_{1:M}}$ $=\begin{bmatrix}h^{11}&\ldots&h^{1M}\\ h^{21}&\ldots&h^{2M}\\ \vdots&\ddots&\vdots\\ h^{N1}&\ldots&h^{NM}\\ \end{bmatrix}+\begin{bmatrix}\sigma_{z}^{11}&\ldots&\sigma_{z}^{1M}\\ \sigma_{z}^{21}&\ldots&\sigma_{z}^{2M}\\ \vdots&\ddots&\vdots\\ \sigma_{z}^{NM}&\ldots&\sigma_{z}^{nM}\\ \end{bmatrix}\odot\epsilon_{\mathbf{z}}$ (10)
$=\mathbf{H(\bm{\chi}^{T_{1:M}}_{xyz}})+\Omega\odot\sigma_{\mathbf{z}}$ (11)
where $\sigma_{z}^{nm}$ is the standard deviation of the range measurement for radar $n$ and target $m$, as defined by the variance in Equation 4, while $\sigma_{z}$ is standard normal Gaussian PDF. Therefore the final state-space equations for the targets is shown in Equations 13 and 12:
@@
$\mathbf{z}^{T_{1:M}}_{k+1}$ $=\mathbf{H}\left((\bm{\chi}^{T_{1:M}}_{xyz})_{k+1}\right)+\Omega\odot\sigma_{\mathbf{z}}$ (13)
@@ 5.2 Radar Kinematics: Second-Order Unicycle Model
With the radars’ positions defined relative to control inputs, the following section outlines the derivation of the FIMs with respect to the targets’ positions. This is based on the radar-target measurement model outlined in Equations 4 and 10, and the radars’ positions parameterized by Equations 14 and 15.
@@ 6 Fisher Information Matrix
$J(\bm{\chi}^{T_{m}}_{xyz};\bm{\chi}^{R_{n}}_{xyz})$ $=(\bm{\chi}^{T_{m}}_{xyz}-\bm{\chi}^{R_{n}}_{xyz})(\bm{\chi}^{T_{m}}_{xyz}-\bm{\chi}^{R_{n}}_{xyz})^{T}\times$
$\left(\frac{4}{K\lVert\bm{\chi}^{R}_{xyz}-\bm{\chi}^{T}_{xyz}\rVert^{6}}+\frac{8}{\lVert\bm{\chi}^{R}_{xyz}-\bm{\chi}^{T}_{xyz}\rVert^{4}}\right)$ (18)
$J(\bm{\chi}^{T_{m}}_{xyz};\bm{\chi}^{R_{1:N}}_{xyz})$ $=\sum_{n=1}^{N}J(\bm{\chi}^{T_{m}}_{xyz};\bm{\chi}^{R_{n}}_{xyz})$ (19)
$J(\bm{\chi}^{T_{1:m}}_{xyz};\bm{\chi}^{R_{1:n}}_{xyz})=$
$\begin{bmatrix}J(\bm{\chi}^{T_{1}}_{xyz};\bm{\chi}^{R_{1:N}}_{xyz})&\!\!0&\!\!\ldots&\!\!0\\ 0&\!\!J(\bm{\chi}^{T_{2}}_{xyz};\bm{\chi}^{R_{1:N}}_{xyz})&\!\!\ldots&\!\!0\\ \vdots&\!\!\vdots&\!\!\ddots&\!\!\vdots\\ 0&\!\!0&\!\!\ldots&\!\!J(\bm{\chi}^{T_{M}}_{xyz};\bm{\chi}^{R_{1:N}}_{xyz})\end{bmatrix}$ (20)
@@
$J(\bm{\chi}^{T_{n}};\bm{\chi}^{R_{1:N}}_{xyz})=\begin{bmatrix}J(\bm{\chi}^{T_{n}}_{xyz};\bm{\chi}^{R_{1:N}}_{xyz})&0\\ 0&0\end{bmatrix}$ (23)
For discrete time filtering systems, such as Equations 13 and 12, the Equation 22 is recursively computed at time $k+1$ from time step $k$ using the famous recursion updates in [12]. The recursion equations are written as
@@
After substituting equations Equations 28, 29 and 30 into Equation 24, and leveraging the Woodbury matrix identity [26], we arrive at [44]:
$\bm{J}_{k+1}(\bm{\chi}^{T_{1:M}}_{k+1})$ $=(W+A^{T}\bm{J}_{k}^{-1}A)^{-1}+$
$\quad\quad E_{f(\bm{\chi}^{T_{1:M}}_{1:k+1})}\left[-\bm{J}_{\bm{D}_{k}}(\bm{\chi}^{T_{1:M}}_{k+1}))\right]$ (32)

### v3

@@ Abstract
Continuously optimizing sensor placement is essential for precise target localization in various military and civilian applications. While information theory has shown promise in optimizing sensor placement, many studies oversimplify sensor measurement models or neglect dynamic constraints of mobile sensors. To address these challenges, we employ a range measurement model that incorporates radar parameters and radar-target distance, coupled with Model Predictive Path Integral (MPPI) control to manage complex environmental obstacles and dynamic constraints. We compare the proposed approach against stationary radars or simplified range measurement models based on the Root Mean Squared Error (RMSE) of the Cubature Kalman Filter (CKF) estimator for the targets’ state. Additionally, we visualize the evolving geometry of radars and targets over time, highlighting areas of highest measurement information gain, demonstrating the strengths of the approach. The proposed strategy outperforms stationary radars and simplified range measurement models in target localization, achieving a 38-74% reduction in mean RMSE and a 33-79% reduction in the upper tail of the 90% Highest Density Interval (HDI) over 500 Monte Carlo (MC) trials across all time steps. Code available at https://github.com/mlpotter/OptimalTrackingAndControl.
@@ 1 INTRODUCTION
(a) Search and rescue in forest fire
(b) Enemy drone localization and tracking
Figure 1: Subfigure (a) shows the civilian application of a search and rescue mission where mobile radars navigate through a dense forest (on fire) to quickly localize and track lost victims. Subfigure (b) shows the military application of enemy drone localization and tracking in a city environment during a cloudy day.
@@
A key challenge in radar-based target tracking is in incorporating realistic sensor models into the information metrics used for controlling the sensors. In particular, most existing control strategies for this problem use the Posterior Fisher Information Matrix (PFIM) [12] (inverse of the Posterior Cramer Rao Lower Bound (PCRLB)) to quantify the relationship between radar-target geometry and target localization performance [13, 14, 15]. However, studies optimizing sensor placement often oversimplify by using a constant covariance matrix for range measurements, neglecting signal path attenuation effects [4]. The approach in [4] focuses on optimizing the angle subtended between the radars and targets but does not encourage the mobile radars to move closer to the target. In [16], the authors provide the error covariance matrix of a Best Linear Unbiased Estimator (BLUE) estimator for the range measurement using a first-order Taylor Series approximation of a distance dependent range (DDR) measurement model, which yields the Fisher Information Matrix (FIM). Similarly, we calculate the FIM for the DDR measurement model from [16], where the covariance matrix of the range measurement is directly proportional to the received power at the radar after signal path attenuation, but we do not apply a first-order Taylor Series approximation to the DDR measurement model. This strategy encourages mobile radars to move closer to the targets, thereby increasing the Signal Noise Ratio (SNR) and improving radar range measurements.
@@
Another major challenge in scenarios with moving targets and movable radars is that the optimal radar-target geometry may change substantially over time. Consequently, many studies use Model Predictive Controller (MPC) [17] for choosing the control inputs of mobile sensors over a finite horizon in response to uncertain target state estimates [18, 19, 20, 21, 22]. However, natural choices for the MPC objective in this problem, such as minimizing/maximizing the D-optimality of PCRLB/PFIM, lead to challenging nonlinear optimization problems that are computationally prohibitive to solve exactly. To address this issue, we instead utilize Model Predictive Path Integral (MPPI) to approximate the optimal control input distribution for the mobile radars. MPPI enables the handling of discontinuous, non-smooth objectives through highly parallelized computation for real-time control [23, 24]. However, as the number of agents, planning horizon, and the actuation space increases, parameter estimation, particularly for the covariance, requires an exponentially growing number of samples. To mitigate this issue and reduce the number of proposal distribution updates and trajectory samples needed in the high-dimensional control space, we integrate advanced Adaptive Importance Sampling (AIS) techniques, such as cross-entropy weighting with covariance shrinkage [25], to prevent overfitting of the sample covariance matrix.
@@
This paper introduces a improved pipeline for continuously optimizing radar placement, aiming to overcome the previously mentioned challenges. We focus on range-only radars, deriving objective functions using FIMs that incorporate signal path attenuation in the range measurement model. These functions are optimized with respect to radar control inputs using MPPI control. Crucially, MPPI control, which is rapidly gaining popularity in control research, offers enhanced flexibility for real-time optimization of nonlinear, discontinuous, and non-differentiable objectives. These advantages enable MPPI to handle future extensions of the proposed objective, such as incorporating more complex signal measurement models or adding more constraints through occupancy grid maps.
@@ 2 Related Work
We explore the literature on optimizing sensor trajectories using the FIM of the likelihood function of the range measurements, where the likelihood function parameters correspond to the target positions. Various forms of the FIM establish a fundamental estimation lower bound limit for the mean squared error covariance matrix [26]. Therefore, many authors construct an objective as the D/A-optimality [27] of the PFIM, parameterized by sensor and estimated target positions [28]. We divide these works into two categories: those that consider static-sensors versus those that consider dynamic-sensors.
@@ 2.2 Dynamic Sensors
Similar to information-based MPC objectives like the D/A optimality of the FIMs, [37] minimizes the mutual information gain between the target state distributions before and after the update step of a linear Kalman Filter, while parameterizing the measurement noise covariance based on the robot’s state. However, the measurement covariance only scales linearly with the sensor-target distance and is clipped at an extremely small variance of 0.01 meters, restricting experiments to small square arenas of 40 to 60 meters.
@@ 3 Preliminaries
    Perfect data association and target detection. We assume that the process of selecting which measurement(s) to incorporate into the target state estimator, is perfect, and the radars constantly track the targets.
@@
  4. 4.
    We assume targets follow a constant velocity model with acceleration process noise and therefore do not account for drastic target maneuvers in our model.
@@ 4 Radar Signal Model
and the set of orthonormal lowpass equivalent signals $s(t)$ such that there is no signal interference when multiple radars transmit simultaneously. The return signal noise $\epsilon_{a}(t)$ is circularly symmetric, zero-mean, complex Gaussian noise with autocorrelation function $\sigma_{a}^{2}\delta(\tau)$. $\zeta$ is the amplitude of the return signal, where $|\zeta|^{2}$ is the radar power return specified by the radar equation [39]:
@@
We denote all the radar parameters, signal parameters, and noise parameters as a constant $\Gamma=\frac{\sigma_{a}^{2}\pi L}{2P_{t}\Lambda_{t}\Lambda_{r}\Xi}$. For brevity, we define shorthand notation as $\mu_{z^{nm}}=2\lVert\bm{\chi}^{R_{n}}_{xyz}-\bm{\chi}^{T_{m}}_{xyz}\rVert$ and $\sigma_{z^{nm}}=\Gamma\lVert{\bm{\chi}^{R_{n}}_{xyz}-\bm{\chi}^{T_{m}}_{xyz}}\rVert^{4}$.
@@ 5.1 Target Kinematics: Constant Velocity Model
The target kinematic model for a single target follows constant velocity motion with acceleration noise [41, 42], which specifies the single-target transition model:
$\bm{\chi}^{T}_{k+1}$ $=A_{\text{single}}\bm{\chi}_{k}^{T}+\epsilon_{w}$ (5)
@@
$A_{\text{single}}$ $=\begin{bmatrix}\bm{I}_{3\times 3}&\Delta t~{}\bm{I}_{3\times 3}\\ \mathbf{0}_{3\times 3}&\bm{I}_{3\times 3}\end{bmatrix}$ (6)
@@
$\bm{W}_{\text{single}}$ $=\bm{W}_{\Delta t}\mathbf{\Sigma}_{w}\bm{W}_{\Delta t}^{T}$ (7)
where
$\bm{W}_{\Delta t}=\begin{bmatrix}\frac{1}{2}\Delta t^{2}~{}\bm{I}_{3\times 3}\\ ~{}\Delta t~{}~{}\bm{I}_{3\times 3}\end{bmatrix}~{},~{}\Sigma_{w}$ $=\bm{I}_{3x3}~{}\sigma_{W}^{2}$ (8)
and $\sigma_{W}$ determines the degree of deviation of the dynamics from constant velocity via the strength of the acceleration variance. Thus, the targets’ transition model (for all targets concatenated as a column vector) is:
@@
$\mathbf{z}^{T_{1:M}}$ $=\begin{bmatrix}\mu_{z^{11}}&\ldots&\mu_{z^{1M}}\\ \mu_{z^{21}}&\ldots&\mu_{z^{2M}}\\ \vdots&\ddots&\vdots\\ \mu_{z^{N1}}&\ldots&\mu_{z^{NM}}\\ \end{bmatrix}+\begin{bmatrix}\sigma_{z^{11}}&\ldots&\sigma_{z^{1M}}\\ \sigma_{z^{21}}&\ldots&\sigma_{z^{2M}}\\ \vdots&\ddots&\vdots\\ \sigma_{z^{N1}}&\ldots&\sigma_{z^{NM}}\\ \end{bmatrix}\odot\epsilon_{\mathbf{z}}$ (11)
$=\mathbf{H(\bm{\chi}^{T_{1:M}}_{xyz}})+\Omega\odot\epsilon_{\mathbf{z}}$ (12)
where $\sigma_{z^{nm}}$ is the standard deviation of the range measurement for radar $n$ and target $m$, as defined by the variance in Equation 4, while $\epsilon_{\mathbf{z}}$ is standard normal Gaussian PDF. Therefore the final state-space equations for the targets is shown in Equations 14 and 13:
@@
$\mathbf{z}^{T_{1:M}}_{k+1}$ $=\mathbf{H}\left((\bm{\chi}^{T_{1:M}}_{xyz})_{k+1}\right)+\Omega\odot\epsilon_{\mathbf{z}}$ (14)
@@ 5.2 Radar Kinematics: Second-Order Unicycle Model
With the radars’ positions defined relative to control inputs, the following section outlines the derivation of the FIMs with respect to the targets’ positions. This is based on the radar-target measurement model outlined in Equations 4 and 11, and the radars’ positions parameterized by Equations 15 and 16.
@@ 6 Fisher Information Matrix
$\bm{J}(\bm{\chi}^{T_{m}}_{xyz};\bm{\chi}^{R_{n}}_{xyz})$ $=(\bm{\chi}^{T_{m}}_{xyz}-\bm{\chi}^{R_{n}}_{xyz})(\bm{\chi}^{T_{m}}_{xyz}-\bm{\chi}^{R_{n}}_{xyz})^{T}\times$
$\left(\frac{4}{\Gamma\lVert\bm{\chi}^{R}_{xyz}-\bm{\chi}^{T}_{xyz}\rVert^{6}}+\frac{8}{\lVert\bm{\chi}^{R}_{xyz}-\bm{\chi}^{T}_{xyz}\rVert^{4}}\right)$ (19)
$\bm{J}(\bm{\chi}^{T_{m}}_{xyz};\bm{\chi}^{R_{1:N}}_{xyz})$ $=\sum_{n=1}^{N}\bm{J}(\bm{\chi}^{T_{m}}_{xyz};\bm{\chi}^{R_{n}}_{xyz})$ (20)
$\bm{J}(\bm{\chi}^{T_{1:m}}_{xyz};\bm{\chi}^{R_{1:n}}_{xyz})=$
$\begin{bmatrix}\bm{J}(\bm{\chi}^{T_{1}}_{xyz};\bm{\chi}^{R_{1:N}}_{xyz})&\!\!0&\!\!\ldots&\!\!0\\ 0&\!\!\bm{J}(\bm{\chi}^{T_{2}}_{xyz};\bm{\chi}^{R_{1:N}}_{xyz})&\!\!\ldots&\!\!0\\ \vdots&\!\!\vdots&\!\!\ddots&\!\!\vdots\\ 0&\!\!0&\!\!\ldots&\!\!\bm{J}(\bm{\chi}^{T_{M}}_{xyz};\bm{\chi}^{R_{1:N}}_{xyz})\end{bmatrix}$ (21)
@@
$\bm{J}(\bm{\chi}^{T_{n}};\bm{\chi}^{R_{1:N}}_{xyz})=\begin{bmatrix}\bm{J}(\bm{\chi}^{T_{n}}_{xyz};\bm{\chi}^{R_{1:N}}_{xyz})&0\\ 0&0\end{bmatrix}$ (24)
For discrete time filtering systems, such as Equations 14 and 13, the Equation 23 is recursively computed at time $k+1$ from time step $k$ using the famous recursion updates in [12]. The recursion equations are written as
@@
After substituting equations Equations 29, 30 and 31 into Equation 25, and leveraging the Woodbury matrix identity [26], we arrive at [45]:
$\bm{J}_{k+1}(\bm{\chi}^{T_{1:M}}_{k+1})$ $=(\bm{W}+\bm{A}^{T}\bm{J}_{k}^{-1}\bm{A})^{-1}+$
$\quad\quad E_{f(\bm{\chi}^{T_{1:M}}_{1:k+1})}\left[\bm{J}_{\bm{D}_{k}}(\bm{\chi}^{T_{1:M}}_{k+1}))\right]$ (33)